# From Wavefunctional Entanglement to Entangled Wavefunctional Degrees of Freedom

### v2

@@ V A Comparison with Entanglement Swapping
Consider two quantum subsystems, each comprising a pair of initially entangled photons; assume the absence of any direct, physical interactions, or dynamical couplings whatsoever between these subsystems. Now, a suitable kind of projective measurement, such as a joint Bell-state measurement, on two of such photons, each drawn from a distinct initially entangled pair, will project the other two into an entangled state. Moreover, the generation of the results of the Bell-state measurement on two particles heralds that the other two particles have been entangled, thereby allowing, for example, the performance of “event-ready detections” of the entangled particles [103].
Typically, various generalizations of the entanglement swapping scheme entail using pairs of initially entangled photons and projecting the state of two initially unentangled particles onto an entangled state. In contrast, the present scheme transforms the entanglement between optical modes (the wavefunctions) into entanglement between the physical properties of the photons (the wavefunctional degrees of freedom), which are described by these modes. As was pointed out in Ref. [103]: “One could have many different kinds of entanglements to begin with, perform various different measurements, and obtain various kinds of entanglement for the emerging particles.” The present approach, therefore, is a form of entanglement swapping that transforms mode-mode entanglement into particle-particle entanglement.
@@ VI Consequences of Non-Ideal Photodetection Efficiencies
Realistically, the detection screen B can be implemented by a single-photon-sensitive EMCCD (Electron Multiplying Charge-Coupled Device) camera. State-of-the-art EMCCD cameras, typically, have quantum efficiencies (equivalent to the photodetection efficiency, above), that can range up to in the visible wavelengths. Moreover, such cameras have advanced features, such as vacuum thermoelectric cooling and electronic optimization of clock-induced charge, for minimizing spurious photodetection events and the dark noise.
@@
One intriguing idea might to be borrow ideas and concepts from standard quantum error-correction theory to detect and mitigate these ancillary-photon-loss errors. For example, a key finding from recent experiments with neutral atom-based quantum processors is that atom-loss-type leakage errors, if detectable and convertible to erasure errors, are easier to correct than other generic classes of unknown errors [107, 108, 109, 110]. Inspired by these ideas pointing to the feasibility of correcting single-quantum-losses, I wish to propose a quantum error-correcting-based strategy, for mitigating the above-described ancillary-photon-loss errors, that is based on classical control, and that is specific to the experiment in Sec. IV.
@@
Let us imagine that the single-photon source, S is deterministically triggered by a clock signal. The frequency of the clock signal, and, therefore, the single-photon emission rate is kept sufficiently low, such that the time interval between the emissions of two successive photons is greater than the sum of the time it takes the single photon to travel from S to the screen, the time for this photon to be detected and registered, and the time required for a classical, logical AND operation that is described in the next paragraph. Unlike the experiment described in Sec. IV, let us not discard the bottom half of the screen, but rather implement the entire screen with a EMCCD camera that is symmetrically positioned with respect to the dotted horizontal line shown in Fig. 3(a). Let us now equip this camera with two measurement read-out channels, R1 and R2, such that R1 outputs the photocurrent generated due to single-photon-detection events in the top half of the camera screen, and R2 outputs the photocurrent generated due to single-photon-detection events in any part of the camera screen; such a situation can be readily arranged by writing a suitable piece of code with the software that controls the camera. The channel R1 is connected to the amplifier, A and the anharmonic potential controller, thereby providing the probabilistic triggering mechanism required in Fig. 2(b).
@@
In contrast, the channel R2 and the clock signal are inputted to a classical, logical AND circuit, and the output is monitored for a window of time equal to the period of the clock signal. If the output is , then no action is taken; however, if the the output is , then it is assumed that the ancillary photon has been lost, and the entire cycle of quantum operation, which commences with the emission of the ancillary photon and concludes with the formation of , is aborted. A new cycle of quantum operation on a freshly prepared, input, entangled state of modes starts with the next clock cycle.
@@
Importantly, the above classical operation should be carried out only after the formation of the macroscopic superposition of the amplifier and the detector, so as not to interfere with the probabilistic altering of the anharmonicity of the potential. In essence, while the efficiency of conversion from to remains unaltered, the output entangled state of photons becomes significantly error-free, at the cost of reduced speed of operations, which is, fundamentally, determined by the single-photon emission rate.
VII Conclusions and Outlook
@@ Acknowledgements.
I would like to thank Chandra Raman for illuminating discussions on the differences between entangled modes and entangled particles, as well as for kindly providing key references on this topic.VIII Data Availability
@@
Figure 3(b) graphically summarizes all these results. Additionally, the entropies of entanglement are found to have no dependence on the relative orientations of the analyzers in space. In contrast, as shown in Fig. 3(c), a rotation in Hilbert space can modulate the expansion coefficients, and, therefore, the degree of entanglement of a mode-mode entangled state. Recently, an approach based on non-Abelian holonomy has been devised to access and tune the angle of this rotation, in a deterministic fashion in a real-world, laboratory setting [84]. Identical results are obtained in Figs. 3(b) and 3(c), if the Rényi entropies are computed, instead of the von Neumann entropies.
@@ Appendix B A Physical Interpretation of the Unitarily Transformed State
Remarkably, the state in Eq. 1, provides a direct and manifestly nonlocal description of the entangled particle pair throughout the flight of the photons. I am assuming that this description is based on, for example, knowledge of successful initial state preparation. In contrast, in the absence of such knowledge, one could interpret the state in Eq. 2, as encoding reconstructed descriptions of quantum events—or assignments of state vectors—after the quantum measurements have been made; notably, similar ideas have been proposed to explain the observation of delayed-choice entanglement swapping [114, 115] in which photons become entangled, after they have already been registered.
@@
To properly interpret , one should regard the individual, recorded events—for instance, measurements that reveal probability amplitudes, as well as relative phases at the two space-like separated, measurement stations—as more fundamental than the quantum state, itself. In this viewpoint, therefore, the wavefunction is the description that the observers assign to the overall situation or the entire phenomenon—as it were, after all the events have occurred—and is determined by the complete set of all possible experimental arrangements; for example, of all allowable relative orientations of and in Fig. 3(a). Of note, this conception of is in harmony with several aspects of the Copenhagen interpretation of quantum mechanics [116], especially, as emphasized by Zeilinger and co-authors in the past few decades [12, 117, 13, 14]. The interpretation of the results of a recent elegant experiment, clarifying the subjective nature of path information, specifically, the “which-way information” in three-particle interferometry, also suggests the primacy of quantum events [118]. This path information plays a key role in how spatial, “mode-entangled,” or path-entangled superpositions of photons are described.
@@ Appendix C A More Formal Justification for the Equivalence
It is well known that the modes, in quantum mechanics, can be mathematically treated as quantum states. In our approach, we model the spatial optical modes, just like any other bosonic mode, as single-particle wavefunctions. Specifically, we map such modes onto the single-quantum, energy eigenmodes of a simple harmonic oscillator. For example, to model the key term of the input state [see Eq. 3], that is, which corresponds to two distinguishable photons occupying two distinct modes, we make the following mappings:
@@
where the spatial coordinates, and , and the particle labels, and have the same meanings, as before. Notice that the above step is the converse of the one shown in Eq. 15, and assumes that these mappings are invertible, which is usually the case. For simplicity, and to avoid any inadvertent pathologies, I am going to assume that this mapping is one-to-one and onto.

### v3

@@ V A Comparison with Entanglement Swapping
Consider two quantum subsystems, each comprising a pair of initially entangled photons; assume the absence of any direct, physical interactions, or dynamical couplings whatsoever between these subsystems. Now, a suitable kind of projective measurement, such as a joint Bell-state measurement, on two of such photons, each drawn from a distinct initially entangled pair, will project the other two into an entangled state. Moreover, the generation of the results of the Bell-state measurement on two particles heralds that the other two particles have been entangled, thereby allowing, for example, the performance of event-ready detections of the entangled particles [115].
Typically, various generalizations of the entanglement swapping scheme entail using pairs of initially entangled photons and projecting the state of two initially unentangled particles onto an entangled state. In contrast, the present scheme transforms the entanglement between optical modes (the wavefunctions) into entanglement between the physical properties of the photons (the wavefunctional degrees of freedom), which are described by these modes. As was pointed out in Ref. [115]: One could have many different kinds of entanglements to begin with, perform various different measurements, and obtain various kinds of entanglement for the emerging particles. The present approach, therefore, is a form of entanglement swapping that transforms mode-mode entanglement into particle-particle entanglement.
Nonetheless, unlike the above approach, standard entanglement swapping protocols rely on projective measurements without external potentials. The introduction of a controlled anharmonic potential signifies a substantial deviation from this standard framework of entanglement swapping. Conceptually, as I have emphasized above, the key difference between this approach and the usual entanglement swapping protocols is that while the former transforms and projects mode-mode entanglement onto particle-particle entanglement, typically, the latter transforms and projects particle-particle entanglement onto a similar form of entanglement, namely, particle-particle entanglement. Operationally, a controlled anharmonic potential appears to be essential, so as to adiabatically manipulate the wavefunctions of the mode-mode entangled states, and reliably and smoothly transform such states into genuinely entangled states of particles of light. Finally, I wish to point out that while my scheme supplements projective measurement on an ancillary photon with a controlled anharmonic potential, it could certainly be possible to supplement projective measurements with other kinds of quantum circuits and systems that reliably and smoothly alter the modal wavefunctions, so as to generate identical final entangled states. The exploration of such additional schemes and protocols, however, is beyond the scope of this paper.
@@ VI Consequences of Non-Ideal Photodetection Efficiencies
Realistically, the detection screen B can be implemented by a single-photon-sensitive EMCCD (Electron Multiplying Charge-Coupled Device) camera. State-of-the-art EMCCD cameras, typically, have quantum efficiencies (equivalent to the photodetection efficiency, above), that can range up to in the visible wavelengths. Moreover, these cameras have advanced features, such as vacuum thermoelectric cooling and electronic optimization of clock-induced charge, for minimizing spurious photodetection events and the dark noise.
@@
One intriguing idea might be to borrow ideas and concepts from standard quantum error-correction theory, so as to detect and mitigate these ancillary-photon-loss errors. For example, a key finding from recent experiments with neutral atom-based quantum processors is that atom-loss-type leakage errors, if detectable and convertible to erasure errors, are easier to correct than other generic classes of unknown errors [121, 122, 123, 124]. Inspired by these ideas pointing to the feasibility of correcting single-quantum-losses, I wish to propose a quantum error-correcting-based strategy, for mitigating the above-described ancillary-photon-loss errors, that is based on classical control, and that is specific to the experiment in Sec. IV.
@@
Let us imagine that the single-photon source, S is deterministically triggered by a clock signal. The frequency of the clock signal, and, therefore, the single-photon emission rate is kept sufficiently low, such that the time interval between the emissions of two successive photons is greater than the sum of the time it takes the single photon to travel from S to the screen, the time for this photon to be detected and registered, and the time required for a classical, logical AND operation that is described in the next paragraph. Unlike the experiment described in Sec. IV, let us not discard the bottom half of the screen, but rather implement the entire screen with a EMCCD camera that is symmetrically positioned with respect to the dotted horizontal line shown in Fig. 5(a). Let us now equip this camera with two measurement read-out channels, R1 and R2, such that R1 outputs the photocurrent generated due to single-photon-detection events in the top half of the camera screen, and R2 outputs the photocurrent generated due to single-photon-detection events in any part of the camera screen; such a situation can be readily arranged by writing a suitable piece of code with the software that controls the camera. Simply put, for every single-photon-detection event, the software will, based on the information from the two measurement read-out channels, R1 and R2, will record two classical bits of information: the first signifying that a successful detection event has occurred; and the second indicating whether this detection has occurred in the top half of the screen. The channel, R1 is connected to the amplifier, A and the anharmonic potential controller, thereby providing the probabilistic triggering mechanism required in Fig. 3(b).
@@
In contrast, the channel, R2 and the clock signal are inputted to a classical, logical AND circuit, and the output is monitored for a window of time equal to the period of the clock signal. If the output is , then no action is taken; however, if the the output is , then it is assumed that the ancillary photon has been lost, and the entire cycle of quantum operation, which commences with the emission of the ancillary photon and concludes with the formation of , is aborted. A new cycle of quantum operation on a freshly-prepared, input, entangled state of modes starts with the next clock cycle.
@@
Importantly, the above classical operation should be carried out only after the formation of the macroscopic superposition of the amplifier and the detector, so as not to interfere with the probabilistic altering of the anharmonicity of the potential. In essence, while the efficiency of conversion from to remains unaltered, the output entangled state of photons becomes significantly error-free, at the cost of reduced speed of operations, which is, fundamentally, determined by the single-photon-emission rate.
It is important to clarify the practical costs, which are associated with the repeated restarts that are necessitated by the above-described logical circuit abort strategy. With regard to resource overhead, this abort strategy requires that the number of times the initial state, is prepared, is increased by a factor of , as the rate of abortion is . Additionally, this strategy requires adding two equipment—such as a clock signal generator and the classical, logical, AND gate—as well as reconfiguring the EMCCD camera, so as to have two measurement read-out channels. Finally, I will conclude this section by offering a few comments about the increased time complexity. In the absence of the abort strategy, one might argue that the total operation time is , where is the response time of the detector, and and are the time scales of operation of the amplifier and the anharmonic potential controller, respectively. This total operation time increases, as a consequence of introducing the abort strategy, to , where is the timescale of operation of the classical, logical, AND gate. More accurately, this increased operation time can be modeled as —where is the time period of the clock signal, and —since the timescale of formation of the final, error-corrected, entangled state of photons is fundamentally limited by . As mentioned earlier, the clock rate is identical to the single-photon-emission rate.
VII Quantum Pre-measurement followed neither by Decoherence nor by Detection
The purpose of this section is to reconsider and re-analyze the Gedankenexperiment, as described in Sec. IV, using the notion of quantum pre-measurement. A key issue that one inadvertently encounters, while analyzing the said Gedankenexperiment, is that one has to, inexorably, assign quantum mechanical wavefunctions to detectors and amplifiers, which are, usually, classical and macroscopic objects. In this section, I will further simply this Gedankenexperiment to its simplest and most essential elements. The screen and the conventional detector is replaced by a single detector atom, which is treated quantum mechanically and as a two-level quantum system; furthermore, there is no longer any amplifier. In what follows, I will briefly survey the usual theory of quantum pre-measurement followed by decoherence—and/or detection—and subsequently justify the entangling scheme of Sec. IV, by utilizing the concept of a quantum pre-measurement of a quantum system by a microscopic, well-isolated, quantum meter.
Typically, in any theoretical formulation of the process of quantum measurement [125, 126, 127], the Hilbert space of the entire universe is factorized into those of the quantum system, S; the measuring apparatus, A; and the rest of the universe, or the environment, E. The quantum system, S is assumed to be small enough, and/or isolated enough, such that its initial state, or the so-called ready state, can be described as a genuine, coherent, linear superposition of two or more eigenstates of the relevant quantum observable. Usually, A is assumed to have multiple macroscopic, accessible degrees of freedom, whereas E is supposed to contain multiple microscopic, hidden or distant degrees of freedom—corresponding to the rest of the environment—that we are not keeping track of.
First, during a process known as a pre-measurement, A, as a consequence of unitary time evolution, measures S and becomes entangled with S. The value of the pointer observable of A becomes correlated with the corresponding eigenvalue of the observable of S, which A is supposed to measure. Notice that A interacts directly with S. Second, if A is sufficiently macroscopic and/or not sufficiently well-isolated, then it gets entangled with the rest of the environment, in a process known as decoherence or environmental entanglement. Physically, the environment is continuously monitoring the combined system of , and, consequently, A splits into preferred pointer basis states that are robust with respect to continual monitoring by the environment. In equations, this two-step process can be summarized as:
wherein the environmental states become orthogonal: , within a very short timescale, namely, the decoherence timescale, such that the two pieces of the final, decoherent, overall wavefunction cannot interfere with each other. The process of decoherence is, therefore, sometimes described as having branched the overall wavefunction into causally distinct, non-interacting worlds that are decoherent with respect to each other.
Now, with the above discussion in mind, let us consider the simplified Gedankenexperiment, as shown in Fig. 4. The atom, A is assumed to be macroscopically delocalized—that is, not in a definite eigenstate of position—along the diffraction plane, and to behave as a quantum meter, namely, a microscopic, measuring apparatus, which is treated quantum mechanically. Moreover, consider this atom to be well-modeled by a two-level system, such that its ground state corresponds to an unionized atomic state and its excited state corresponds to an ionized atomic state, and assume the difference between the excited and the ground energy eigenstates of this atom to be identical to the energy of photon 3. Notice that there are no longer any amplifiers or macroscopic detectors.
As has been discussed in Sec. IV, the ancillary photon 3 can be described by the state: , as it transits through the double-slit apparatus. In accordance with the analysis of this double-slit diffraction experiment, in the textbook of Schiff [111], the overall state of the macroscopically delocalized detector atom and the single photon can be described as:
where and are the two energy eigenstates of the detector atom, and the expression for the probability amplitude, —specifically, the atomic transition probability amplitude—is given in Ref. [111]. Notice that the detector atom acts as a quantum meter and that the total excitation number, of this atom, is the salient pointer observable. Subsequent to the formation of the above superposition of an excited atom—with an accompanying ejected photo-electron—and a non-excited atom—without any ejected photo-electron—a quantum mechanical photocurrent is produced that corresponds to a superposition of the form: . This quantum mechanical photocurrent probabilistically triggers the anharmonic potential controller; a value of is supposed to indicate triggering. Crucially, I am assuming that the anharmonic potential controller is spatially located sufficiently close to the detector atom, such that the anharmonic potential controller is within the region of successful capture of the photo-electron; furthermore, it is assumed that the anharmonic potential controller is suitably grounded, so that a closed circuit is established for the flow of photo-electrons.
It is essential to emphasize that both the detector atom and the single-electron-based triggering mechanism—that can be realized, for instance, using single-electron turnstile devices [128, 129]— of the anharmonic potential controller are all assumed to be microscopic, quantum mechanical, and sufficiently well-isolated from the rest of the environment, so that there is neither any decoherence nor any detection nor any collapse of the above superpositions. From this point onward, the remainder of the analysis of how the final entangled state of photons is formed is identical to what has been described in Sec. IV, specifically, from the presentation of the Hamiltonian in Eq. (5) onward.
VIII Conclusions and Outlook
@@ Acknowledgements.
I would like to thank Sara Sloman, J. Van Butcher, Chandra Raman, and Xueda Wen, for many delightful and helpful discussions—on related topics—that inspired me to work on this problem. Additionally, I wish to thank Rajesh Gopakumar and Abhishek Dhar, for hosting me at the ICTS–TIFR, Bengaluru, where a part of this paper was written.IX Data Availability
@@
Figure 5(b) graphically summarizes all these results. Additionally, the entropies of entanglement are found to have no dependence on the relative orientations of the analyzers in space. In contrast, as shown in Fig. 5(c), a rotation in Hilbert space can modulate the expansion coefficients, and, therefore, the degree of entanglement of a mode-mode entangled state. Recently, an approach based on non-Abelian holonomy has been devised to access and tune the angle of this rotation, in a deterministic fashion in a real-world, laboratory setting [86]. Identical results are obtained in Figs. 5(b) and 5(c), if the Rényi entropies are computed, instead of the von Neumann entropies.
@@ Appendix B A Physical Interpretation of the Unitarily Transformed State
Remarkably, the state in Eq. (1), provides a direct and manifestly nonlocal description of the entangled particle pair throughout the flight of the photons. I am assuming that this description is based on, for example, knowledge of successful initial state preparation. In contrast, in the absence of such knowledge, one could interpret the state in Eq. (2), as encoding reconstructed descriptions of quantum events—or assignments of state vectors—after the quantum measurements have been made; notably, similar ideas have been proposed to explain the observation of delayed-choice entanglement swapping [133, 134] in which photons become entangled, after they have already been registered.
@@
To properly interpret , one should regard the individual, recorded events—for instance, measurements that reveal probability amplitudes, as well as relative phases at the two space-like separated, measurement stations—as more fundamental than the quantum state, itself. In this viewpoint, therefore, the wavefunction is the description that the observers assign to the overall situation or the entire phenomenon—as it were, after all the events have occurred—and is determined by the complete set of all possible experimental arrangements; for example, of all allowable relative orientations of and in Fig. 5(a). Of note, this conception of is in harmony with several aspects of the Copenhagen interpretation of quantum mechanics [135], especially, as emphasized by Zeilinger and co-authors in the past few decades [12, 136, 13, 14]. The interpretation of the results of a recent elegant experiment, clarifying the subjective nature of path information, specifically, the which-way information in three-particle interferometry, also suggests the primacy of quantum events [137]. This path information plays a key role in how spatial, mode-entangled, or path-entangled superpositions of photons are described.
@@ Appendix C A More Formal Justification for the Equivalence
It is well known that the modes, in quantum mechanics, can be mathematically treated as quantum states. In our approach, we model the spatial optical modes, just like any other bosonic mode, as single-particle wavefunctions. Specifically, we map such modes onto the single-quantum, energy eigenmodes of a simple harmonic oscillator. For example, to model the key term of the input state [see Eq. (3)], that is, which corresponds to two distinguishable photons occupying two distinct modes, we make the following mappings:
@@
where the spatial coordinates, and , and the particle labels, and have the same meanings, as before. Notice that the above step is the converse of the one shown in Eq. (19), and assumes that these mappings are invertible, which is usually the case. For simplicity, and to avoid any inadvertent pathologies, I am going to assume that this mapping is one-to-one and onto.